OpenStudy (anonymous):

If f(x) = 2 x^{4 x}, find f '( 2 ).

hartnn (hartnn):

could you find the derivative f'(x) first ?

OpenStudy (anonymous):

that's the thing I tried, but its wrong

hartnn (hartnn):

can you show your work ? i'll try and find the error

OpenStudy (anonymous):

2*(4x)x^(4x-1) this isn't right thought

OpenStudy (anonymous):

I also tries chain rule. but what would be g(x)??

hartnn (hartnn):

do u know logarithmic differentiation ?

OpenStudy (anonymous):

\[\Large f(x)=2x^{4x}\] or can be written as \[\Large y=2x^{4x}\] take natural log of both sides \[\Large \ln(y)=\ln(2x^{(4x)})\]

OpenStudy (anonymous):

now expand the right side little bit using property of log \[\Large \ln(y) =\ln(2) +\ln(x^{4x})\] \[\Large \ln(y)=\ln(2)+4xln(x)\] differentiating both sides with respect to x \[\Large \frac{d(\ln(y)}{dx}=\frac{d(\ln(2)}{dx}+\frac{d}{dx} (4x*(\ln(x))\] since derivative of constant is zero so it becomes \[\Large \frac{d(\ln(y)}{dx}=\frac{d}{dx} (4x*(\ln(x))\] just apply product rule on right side. let me know result after applying product rule.

OpenStudy (anonymous):

ok but wouldn't the left side be ln'(y)

OpenStudy (anonymous):

it will be but \[\Large \ln'(y)=\frac{1}{y}\frac{dy}{dx}\]

OpenStudy (anonymous):

for right side: 4xln(x) + 4

OpenStudy (anonymous):

thats correct

OpenStudy (anonymous):

the left side as i wrote above is \[\Large \frac{1}{y}\frac{dy}{dx}\]

OpenStudy (anonymous):

writing both together \[\Large \frac{1}{y}\frac{dy}{dx}=4\ln(x)+4\]

OpenStudy (anonymous):

but \[\Large y=2x^{4x}\]

OpenStudy (anonymous):

isn't it 4xln(x) not 4ln(x)

OpenStudy (anonymous):

no it should be 4ln(x)+4 after product rule because \[\Large \frac{d}{dx}(4x)=4\]

OpenStudy (anonymous):

let me write product rule expansion \[\Large \frac{d}{dx}(4x*\ln(x)=\frac{d}{dx}(4x)\ln(x)+ 4x \frac{d}{dx}(\ln(x))\]

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

going back to previous result \[\Large \frac{dy}{dx}=y(4\ln(x+4)\] but \[\Large y=2x^{4x}\] replace the value of y in the above expression. let me know the result.

OpenStudy (anonymous):

@jolee did you get it ?

OpenStudy (anonymous):

im supposed to find y' though not d/dx

OpenStudy (anonymous):

put u put it all on one page

OpenStudy (anonymous):

please help fast!!!

OpenStudy (anonymous):

these all notations are same. \[\large \frac{dy}{dx}=y'=f(x)\]

OpenStudy (anonymous):

yes but f'(x)

OpenStudy (anonymous):

\[\large \frac{dy}{dx}=y'=f'(x)\]

OpenStudy (anonymous):

after replacement of y you should have \[\Large \frac{dy}{dx}=f'(x)=2x^{4x}(4\ln(x)+4)\]

OpenStudy (anonymous):

just substitute x=2 in the above expression and use calculator to get f'(2) .

OpenStudy (anonymous):

are u certain this is correct?

OpenStudy (anonymous):

yes it is . you can simply it little more if you want.

OpenStudy (anonymous):

is the answer 181708?

OpenStudy (anonymous):

actually 443848

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no did you convert your calculator to radians ?

OpenStudy (anonymous):

yes im in radian mode. is this wrong?

OpenStudy (anonymous):

why would that matter?

OpenStudy (anonymous):

all calculations in calculus are done in radians. try again because i am getting 3467.566

OpenStudy (anonymous):

but why in radians? srry.

OpenStudy (anonymous):

just for the sake of convience.

OpenStudy (anonymous):

I keep getting 443848. first from ln part, 6.77 times 4^8

OpenStudy (anonymous):

2*2^8*6.77

OpenStudy (anonymous):

ahh thank you. do u also know physics by chance?

OpenStudy (anonymous):

by chance i do know. :)

OpenStudy (anonymous):

ill ask a question in physics section then!